What is the output of the following program? Show each step  to get to the output

#include <stdio.h>

int f(int* a, int* b);

int main() {
       int a = 6, b = 7, c = 8;

       b = f(&c, &a);

       printf("a = %d, b = %d, c = %d\n", a, b, c);   

       return 0;

}

int f(int* a, int* b) {
       *a = *a + *b;
       *b = *a - *b;
       printf("a = %d, b = %d\n", *a, *b);
       return 3*(*b);

}
Answer #1
b = f(&c, &a);
f(8,6)
Here the function call is b = f(&c, &a);
and the function definition is int f(int* a, int* b)
Here, variable c in main represents variable a in function f
and variable a in main represents variable b in function f

a will be 8 and b will be 6
*a = *a + *b = 8 + 6 = 14
*b = *a - *b = 14 - 6 = 8
Then it prints
a = 14, b = 8
and then returns 3*8 = 24
Then main prints 
a = 8, b = 24, c = 14

So, the output is
a = 14, b = 8
a = 8, b = 24, c = 14